Y = 2^3x + 1/2 ln x + 4 t^3
1
Expert's answer
2020-03-03T07:56:07-0500

Y = 2^3x + 1/2 ln x + 4 x^3

Solution:-

Y="(2^3)^x" "=8^x"

Y="8^x+{1\above{2pt} 2}lnx+4x^3"

Graph

"Y=8^x+{1\above{2pt} 2}lnx+4x^3"